The length of a rectangular flower bed is 11 ft. less than 3 times its width. The area of the bed is 20ft^2. Find the dimensions of the flower bed.
Answer by mananth(16949)

The length of a rectangular flower bed is 11 ft. less than 3 times its width. The area of the bed is 20ft^2. Find the dimensions of the flower bed.
let width be x
three times width =3x
length = 3x-11
x(3x-11)=20
3x^2-11x-20=0
3x^2-15x+4x-20=0
3x(x-5)+4(x-5)=0
(3x+4)(x-5)=0
x=5 width
width is 5 feet
length = 3x-11
=3*5 -11
=4 feet length